nightshade0773:

To kick off the summer, Maura is planning a "School's Out!" party for all her friends. She and her mom pick up a giant sub and a vat of root beer from Subs & Suds. When they get to the park, they cut the sub into 12 equal sections. Each section is 1/4 of a foot long.

nightshade0773:

Solve this equation for g to find the length of the giant sub.

mhanifa:

Do you have an idea what to start with?

nightshade0773:

no i dont

mhanifa:

Let the length be g, you cut it by 12 equal sections and each section is 1/4 of a foot

mhanifa:

Can you make an equation to reflect this?

harper14:

the answer will be 48 idk if this is right but i gt it by doing 1/4 divided by 12 which gave me 48 again idk if this is right so if its not don't come at meh

mhanifa:

@harper14 wrote:
the answer will be 48 idk if this is right but i gt it by doing 1/4 divided by 12 which gave me 48 again idk if this is right so if its not don't come at meh
This is incorrect

harper14:

well it asked what was is equally eaning have to divided

harper14:

but ok do what you wnna do

mhanifa:

@nightshade0773 can you set an equation? It will be in the form of g / a = b

mhanifa:

g is your total length, a- number of sections, b- length of each section

nightshade0773:

one of the answer choices are g/12=1/4

mhanifa:

This is correct

mhanifa:

Now you need to solve it for g, can you?

nightshade0773:

i got 3

mhanifa:

That's correct. I assume you did: g = 12*1/4 g = 3 feet

mhanifa:

@harper14 wrote:
but ok do what you wnna do
Are you happy with the answer? Did you learn something?

harper14:

no i didn't u weren't very clear but kk good job